OpenStudy (anonymous):

The price of three different size detergents are consecutive odd integers. The sum of the prices is $45. What are the prices of the three detergents? A. $14, $16, $18 B. $13.67, $15.67, $17.67 C. $13, $14, $15 D. $13, $15, $17

OpenStudy (anonymous):

D. $13, $15, $17

OpenStudy (anonymous):

Only Answer D applies because its the only answer that contains consecutive odd integers

OpenStudy (anonymous):

consecutive odd integers means odd whole numbers in a row

OpenStudy (anonymous):

i thought so but wasnt sure i thought it was basically asking for the odd numbers

OpenStudy (anonymous):

. The perimeter of a rectangle is 60 feet. The length is 6 feet more than the width. Find the length and width. A. w = 12, l = 18 B. w = 12, l = 6 C. w = 13.5, l = 19.5 D. w = 27, l = 33

OpenStudy (anonymous):

this ibe i have no idea

OpenStudy (anonymous):

*one

OpenStudy (anonymous):

ok so perimeter is the addition of all the sides

OpenStudy (anonymous):

A. w = 12, l = 18

OpenStudy (anonymous):

P = l + w + l + w = 2 × l + 2 × w

OpenStudy (anonymous):

sorry its taking so long im eating and doing this

OpenStudy (anonymous):

added up 18 *4 and it said 72 i do believe

OpenStudy (anonymous):

im prob doing all wrong though

OpenStudy (anonymous):

im trying i really am sorry

OpenStudy (anonymous):